OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=3/2x y=-1/2+4 (____,_____) i will fan who ever helps me out

RhondaSommer (rhondasommer):

I can help :)

OpenStudy (anonymous):

@K12girlohva who ever can help :)

RhondaSommer (rhondasommer):

because y= -1/2 + 4 replace the y in the first equation

RhondaSommer (rhondasommer):

so now -1/2 +4= 3/2x

RhondaSommer (rhondasommer):

3.5 = 3/2x

RhondaSommer (rhondasommer):

divide 3/2 on both sides and that is your answer :)

OpenStudy (anonymous):

ight thank you so much

RhondaSommer (rhondasommer):

no problemo :)

OpenStudy (anonymous):

so would it be .58 ? @RhondaSommer

RhondaSommer (rhondasommer):

as i am looking at this....is the second equation supposed to be -1.2 x? or just -1/2

OpenStudy (anonymous):

-1/2

RhondaSommer (rhondasommer):

then i am pretty sure its 2 1/3

OpenStudy (anonymous):

o ok thanks

RhondaSommer (rhondasommer):

no prob. pls medal or fan :)

OpenStudy (anonymous):

i already faned ya

RhondaSommer (rhondasommer):

thnx